OpenStudy (anonymous):

WILL MEDAL AND FAN Evan is painting a stool that has a cylindrical seat and four rectangular prism legs. The seat has a diameter of fourteen inches and a height of 4 inches. Each leg is 4 inches by 4 inches by eighteen inches. How much area will Evan have to paint? Use three point one four for pi. Can you please show the work? I have like 5 minutes to get this last question in

OpenStudy (anonymous):

Ok. A=2πrh+2πr2 So with this formula, plug it all in. 2*3.14*7*4 = 175.84 now, we do 2*3.14*7*7 = 307.72 Add them together 175.84+307.72 = 483.56

OpenStudy (anonymous):

Thank you so much, just in time to

OpenStudy (anonymous):

No problem :D